Question 1 9y + 30y + 25y = 0 O y = (C1 + c2x)e-5x/3 O y = (c1 + c2x)25x/3 O y=ce-5x/3 Oy- cze-5x/3 + c2
Answer #1

Giver, find its 9y+ Boy+ 25y = 0 To solve this first we will auxilary eq → 9m2+ 30m + 25 = 0 (3m + 5)² = 0 m = کابی Now, in